Fermi National Laboratory


Accelerator Update

Wednesday January 14
The day shift began with the Tevatron (Tev) in shot setup.

At 10:24 AM a Tev quench occurred at E48U. The Tev had just started to ramp to 980 Gev.

Antiproton Source (Pbar) returned to stacking at 11:01 AM.

CDF and D0 both made accesses into their respective collision halls.

After quench recovery and experiment access the Tev is ramping once again at 1:39 PM. The Tev will be doing studies to determine the cause of the quench and to give Pbar time to increase the stack.

At 11:19 PM the Tev quenched while aborting studies beam. The quench occurred at BA and BB.

Thursday January 15
The midnight shift began with the Tev in quench recovery, Pbar stacking, and with MiniBoone taking beam.

At 1:49 AM the quench recovery is complete and the Tev is ramping once again.

The Tev goes into shot setup at 11:56 AM.

At 2:12 PM, Operations established store #3172 with an initial luminosity of 29.72 E30.

Pbar returns to stacking at 2:44 PM.

Friday January 16
The Midnight shift began with store #3172 colliding, stacking, and with beam to MiniBoone.

MiniBoone turned off at 4:00 AM in order to allow for sufficient cooldown time for morning access in MI-12 B.

Today's Plans
Continue to stack and establish stores

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/16/2004   email Fermilab